Estimate the electric field inside a metalic sphere of radius 10cm and electric charge of -9x10-6C. Explain your answer,
Answer #1

electric field = kQ/R^2

where k = 1/4pi*eo = 9*10^9

Q = charge , R = radius

EF = 9*10^9*9*10-6 / 0.1^2

= 8.1*10^6 N /C